The sidecap package

First, a standard L^{A}TEX figure environment (fig. 1). Since width of the body of the figure is considerably less than the document’s \textwidth, there is much space wasted.

[figure]

Fig. 1: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

For such cases, the sidecap package provides the SCfigure and SCtable environments to put the caption beside the body of the floating environment.

In this document, we used the defaults of the sidecap package:

  \usepackage{sidecap}

In the text, we used the optional arguments for the following figures (fig. 2, 3, and 4).

  \begin{SCfigure} ...
  \begin{SCfigure}[1.2][bhp] ...
  \begin{SCfigure}[][bhp] ...

In the first case, no optional arguments were given, so the defaults ([1.0][tbp]) are used. I.e., the caption will be as wide as the body of the float; and the ‘float placement specifier’ is the same as for L^{A}TEX’s standard floats (fig. 2). In the second case, both optional arguments are given to change the caption width as well as the placement options (fig. 3). To change only the placement, just leave the first optional argument empty (fig. 4).

[figure]

Fig. 2: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]

Fig. 3: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]

Fig. 4: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Now, a standard L^{A}TEX table environment (tab. 1).

Tab. 1: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

1 2 3 4 5 6
2 4 6 8 10 12
3 6 9 12 15 18
4 8 12 16 20 24
5 10 15 20 25 30
6 12 18 24 30 36
7 14 21 28 35 42
8 16 24 32 40 48
9 18 27 36 45 54
10 20 30 40 50 60
11 22 33 44 55 66
12 24 36 48 60 72
13 26 39 52 65 78
14 28 42 56 70 84

The only difference between the SCfigure and the SCtable environments is the vertical position of the caption with respect to the body of the float: The former places the caption bottom-aligned, the latter top-aligned (compare fig. 2 to tab. 2).

Also, like standard L^{A}TEX, the sidecap package doesn’t care about the contents of the float (cf. tabs. 1 and 2).

If you want to use all the remaining space for the caption, just specify

[figure]

Tab. 2: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]

Tab. 3: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

a sufficiently large factor in the first optional argument. E.g., for tab. 3 we used the following:

  \begin{SCtable}[50] ...

Note, however, that TEX may complain

  ! Dimension too large.

if the factor times the width of the body is huge enough to exceed the range TEX is able to represent in a length register. (Max. \maxdimen = 16383.99998pt ≈ 5.76 meters).